A landscape architect designed a decorative grass border around a large community garden. The community garden is 20 feet wide and 40 feet long. The area of the grass border is to be 544 square feet and uniform width.
How wide should the architect make the border?

Answer by scott8148(6628)

(20 + 2w)((40 + 2w) - (20)(40) = 544

800 + 120w + 4w^2 - 800 = 544

w^2 + 30w - 136 = 0

(w - 4)(w + 34) = 0

negative value not realistic ___ w = 4